Circular No. 7565

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001K IN IC 677 Y. L. Qiu and J. Y. Hu, on behalf of the Beijing Astronomical Ob-

servatory (BAO) Supernova Survey, report the discovery of an apparent supernova (mag about 17.0) on unfiltered images taken on Jan. 15.9 UT. SN 2001K is located α = 11^{h}13^{m}56^{s}.06, δ = +12^{o}18′07′′.8 (equinox 2000.0), which is 6′′.7 west and 3′′.0 north of the nucleus of IC 677. W. D. Li, University of California at Berkeley, reports that SN 2001K appears at mag ∼ 16.5 on an unfiltered image taken with the Katzman Automatic Imaging Telescope (KAIT) on Jan. 17.4, from which Li measured the object’s position end figures to be 56^{s}.09, 06′′.6. Nothing appears at this position on a BAO image of the same field taken on 2000 Dec. 20 (limiting mag 18.3) or on a KAIT image taken on Dec. 30.5.

V445 PUPPIS W. Liller, Viña del Mar, Chile, reports that inspection of 54 films

taken from 1982 Oct. 7 to 1985 Nov. 25 (limiting red-orange mag 10–11) reveals no outburst of V445 Pup. He adds that, observing with a CCD and broadband V filter over a period of 6.5 hr (exposures every 5 min) beginning Jan. 14.115 UT, no detectable variation (rms ± 0.07 mag) was found in the brightness of V445 Pup (comparison star GSC 6543.2917, located 2′.7 to the south); neither did a period-search program reveal any significant result. G. Masi, Ceccano, Italy, also reports that V445 Pup remained around mag ∼ 8.7 (unfiltered CCD) between Dec. 31.7 and Jan. 7.0, and that short-term fluctuations appeared to be absent.

COMETS C/1998 U6, C/1998 V5, C/1998 V6, C/1998 X11 (SOHO)

Further to IAUC 7562, D. Hammer reports his measurements for additional Kreutz sungrazing comets found by M. Oates (Dec. comet) and X. Leprette (remaining comets) in C2 data on SOHO website images. The reduced observations and orbital elements by B. G. Marsden are given on the MPECs cited below.

Comet  1998 UT  α_{2000}  δ_{2000}  MPEC
C/1998 U6 Oct. 29.263  14^{h}05.^{m}5  −14^{o}33′  2001-A47
C/1998 V5 Nov. 2.643  14 24.0  −16 01  2001-A43
C/1998 V6  2.643  14 24.9  −15 55  2001-A43
C/1998 X11 Dec. 8.795  17 04.9  −24 32  2001-A43
2001 January 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green